OpenStudy (rebecca1588):

Fan + Medal Maya has m baseball cards. She puts half of her cards in a pile to give to her brother and sister. She gives & 12 of these cards to her brother and then gives the rest of the pile to her sister Trisha. The expression m/2 −12 can be used to find the number of cards Trisha gets from Maya. Maya thinks she started with 20 cards. If Maya is right, then how many cards would Trisha get? ___________

OpenStudy (ayubie):

If Maya thinks she started with 20 cards you can plug in 20 for m. This will give you: 20/2 -12 = 10-12 = -2 The number is negative so Maya is probably wrong, because in this case Trisha will have to give Maya 2 cards.